OpenStudy (anonymous):

Need help! Steven is four years older than Kuldeep. Five years from now, the sum of their ages will be 100. How old is each person right now?

OpenStudy (***[isuru]***):

in the present if Steven is = "x" years old then Kuldeep = "x -4" years old ( since stevan is 4 years older than kuldeep ) but after 5 years which mean Stevan = x + 5 Kuldeep = (x - 4) + 5 = x +1 their sum is 100 ... so........... x + 5 + x + 1 = 100 now solve this for x .. u will get the Stevans age... Subtract 4 from it . u will get Kuldeep age.. can u do it :) ?

OpenStudy (anonymous):

yep i got it thank you :)

OpenStudy (***[isuru]***):

u r welcome :)